Sol Summary – November 4th [status draft

Sol: 6
Summary Title: Embracing fresh opportunities.
Author’s name: David Mateus
Mission Status: Nominal
Sol Activity Summary: The day started at 8 a.m. as the crew gathered for a hearty breakfast, fostering camaraderie through casual conversations. Post-breakfast, the crew’s focus shifted to meticulous planning for tomorrow’s Extravehicular Activities (EVAs), delving into the intricacies of each mission’s objectives and logistics.
The team’s collaborative discussions during the EVA planning session not only enhanced mission preparedness but also deepened their understanding of one another’s strengths and perspectives. This talk provided valuable insights into each crew member’s unique qualities and experiences, reinforcing their cohesion as a unit.
As the day progressed into the afternoon, individual pursuits took center stage. The Crew Astronomer began an exciting endeavor, capturing his first-ever solar images. Meanwhile, the Greenhab Officer diligently continued their work on maintaining and improving the Greenhab, implementing suggestions and guidance provided by the Mission Support team.
This productive day showcased the crew’s versatility and dedication as they seamlessly transitioned between collaborative planning and pursuing their projects, contributing to the overall success of our mission.
Look Ahead Plan: Perform two EVAs and continue with the projects
Anomalies in work: None
Weather: Partly cloudy
Crew Physical Status: Nominal.
EVA: None
Reports to be filed: Sol Summary, Journalist report, Photos, Operations report, Green Hab report, Astronomy report, EVA #5 Request, EVA #6 Request.
Support Requested: Pick up the trash.

Mid-mission Research Report – November 4th

Project 1: The sensor kit was successfully placed in the generator during an EVA. This allowed the engineer to evaluate their fine motor skills while wearing the full EVA suit. Since then, the sensors have been collecting data every night, which so far has been used mainly for debugging the data collection and analysis scripts.

Project 2: The script to recognize the drone image within the larger satellite image is currently in progress, and tests using real data have already started. The real data comes from two EVAs of the Candor Chasma. Having data from two EVAs provide different light conditions on the images to test the robustness of the algorithms. In the next few days the crew will collect more datasets from different areas in order to finish fine tuning the algorithm.

Project 3: A drone will be used for search and rescue, along with specific signs of communication (indications), tomorrow on an EVA in el Dorado Canyon

Project 4: The 3D mapping project using drones and the generation of digital pathways based on them is operating effectively and was successfully tested during EVA 3 and 4.

Project 5: nothing to report

Project 6: For the Mars construction materials project, We have been working with various soil samples collected from strategic locations in the area. Initially, rice-based starch was prepared within the ScienceDome and later combined with the three sieved soil samples to produce three concrete bricks with distinct properties. The three brick units were crafted through the amalgamation of 100 grams of soil and 6 grams of starch, followed by a 4-hour baking process to enhance their structural strength. Subsequently, strength tests will be conducted on the three Martian concrete pieces

Project 7: The Deep Space Observatory project is proving to be successful, capturing various objects daily and testing proposed processing methods to emulate different palettes that typically include the oxygen 3 band, which the observatory’s telescope doesn’t encompass. It aims to test processes that increase the signal-to-noise ratio. The solar observation project commenced today, experimenting with processing methods to extract the halo around solar flares from the chromosphere

Astronomy Report – November 4th



Name: Luis Diaz
Crew: 285
Date: 11/04/2023
Robotic Telescope Requested: MDRS-WF
Objects to be Imaged this Evening: Accumulate more light for Helix Nebula, Horsehead Nebula and maybe others.
Images submitted with this report: –
Problems Encountered: –
Solar Features Observed: Sunspot, very defined filaments and solar flares.
Images submitted with this report: – Full sun (mosaic of 2 photographs)
Problems Encountered: –

Journalist Report – November 3rd

by Marina Buqueras Rovira


Sol 5

7 am Temprano en Marte, el Commander encabeza el despertar de la tripulación. Hoy nos espera una larga actividad extravehicular “EVA”. Chocolate con leche, zumo de naranja, galletas o cereales es nuestro desayuno.

9 am Andrea, David y Luis inician la misión EVA3, la importancia de su salida es conseguir capturar imágenes con el drone para así poder generar un ortomosaico en 3D para la EVA siguiente. Estuvieron casi 3 horas en pleno sol caminando por Candor Chasma, de vuelta Opportuniy y Perseverance, les regresan a la base cansados aunque entusiasmados!

Mientras la tripulación de la EVA3 examinaba la zona y cumplía con su misión, el resto preparaban el almuerzo con cariño.

12 pm Hoy tenemos arroz con lentejas y bacon, también espinacas con champiñones. Unos recuperan la fuerza y los demás cargan las pilas para la EVA4.

13:30 pm Tomás, Andrés y Marina, están listos para su misión. La eficacia de Luis hace posible que nuestra ruta estuviera perfectamente trazada en el mapa, también nos proporcionó información del cañón en 3D para prevenir cualquier peligro. Nuestra misión se enfocó en llegar a Candor Chasma con los rovers Curiosity y Spirit y seguir andando por el cañón. Después de 4 horas llegamos a la base con muestras y con la certeza de haber concluido satisfactoriamente la EVA4.

17 pm La tarde se presenta tranquila, entre bromas y chanzas. Nos preparamos para la CapCom y hoy tiene pinta de noche de juegos!

¡Nos vemos mañana!


Sol 5

7 am Early in the morning on Mars, the Commander leads the crew’s awakening. Today, a long extravehicular activity "EVA" awaits us. Milk chocolate, orange juice, cookies, and cereals make up our breakfast.

9 am Andrea, David, and Luis are beginning the EVA3 mission. The importance of their outing is to capture images with the drone in order to generate a 3D orthomosaic for the next EVA. They spent nearly 3 hours walking under the blazing sun in Candor Chasma. On their return, Opportunity and Perseverance welcomed them back to the base, tired but excited!

While the Crew the EVA3 was examining the area and fulfilling its mission, the rest lovingly prepared lunch.

12 pm Today, we have rice with lentils and bacon, as well as spinach with mushrooms. Some regain their strength, while the others recharge for EVA4.

1:30 pm Tomas, Andres, and Marina are ready for their mission. Luis’s efficiency ensured that our route was perfectly plotted on the map. He also provided us with 3D canyon information to prevent any danger. Our mission focused on reaching Candor Chasma with the Curiosity and Spirit rovers and continuing to walk through the canyon. After 4 hours, we returned to the base with samples and the certainty of having successfully completed EVA4.

5 pm The afternoon presents itself calmly, with jokes and banter. We are preparing for CapCom, and tonight looks like game night!
See you tomorrow!

Sol Summary – November 3rd

Sol: 5
Summary Title: Exploring Martian Labyrinths.
Author’s name: David Mateus
Mission Status: Nominal
Sol Activity Summary: The crew started the day at 7 a.m. to have breakfast and prepare the first. At 9 a.m., the first team left for Candor Chasma; this team overflew Candor Chasma to take images, create a 3D model for future explorations, and take data for some projects. The first EVA team arrived at noon, the whole crew ate lunch and started preparing the second team that left at 1:30 p.m. to Candor Chasma.
The second team entered the canyon following maps and 3D images created by the first team.
Look Ahead Plan: Prepare future EVAs and continue with the projects
Anomalies in work: None
Weather: Partly cloudy
Crew Physical Status: Nominal.
EVA: Two EVAs to Candor Chasma
Reports to be filed: Sol Summary, Journalist report, Photos, Operations report, Green Hab report, Astronomy report, EVA #3 Report, EVA #4 Report.
Support Requested: None

Operations Report – November 3rd

Crew 285 Operations Report 03-11-2023
SOL: 5
Name of person filing report: Tomás Ignacio Burroni
Non-nominal systems: None
Notes on non-nominal systems: N/A
Spirit rover used: Yes
Hours: 233.4
Beginning charge: 100%
Ending charge: 47%
Currently charging: Yes
Opportunity rover used: Yes
Hours: 143.1
Beginning charge: 100%
Ending charge: 57%
Currently charging: Yes
Curiosity rover used: Yes
Hours: 256.4
Beginning charge: 100%
Ending charge: 48%
Currently charging: Yes
Perseverance rover used: Yes
Hours: 275.3
Beginning charge: 99%
Ending charge: 46%
Currently charging: Yes
General notes on rovers: N/A
Summary of Hab operations:
WATER USE: 20 gallons
Water (static tank): 367 gallons
Static tank pipe heater (on or off): on
Static tank heater (On or off): off
Toilet tank emptied: no
Summary of internet: No comments
Summary of suits and radios: Attached is a picture of the marks found on the suit used by the actor yesterday. We are not sure which of these were already present as we had not used the one-piece suits previously.
Summary of GreenHab operations: sowing several plants and watering
WATER USE: 6.0 gallons
Heater: On
Supplemental light: Off
Harvest: None
Summary of ScienceDome operations: Not used today
Dual split: Off
Summary of RAM operations: Not used today
Summary of any observatory issues: None
Summary of health and safety issues: None
Questions, concerns and requests to Mission Support: None

GreenHab Report – November 3rd

GreenHab Officer: Andres David Reina
Environmental control: (heater, fan, door open?)
Open door: 9:30 am:
-     Temperature: 22,11C
-     Humidity:12%
Close door: 18:05
-     Temperature: 23,11C
-     Humidity: 11%

Hours of supplemental light: None
Daily water usage for crops: 6 gallons.
Daily water usage for research and/or other purposes: None
Water in Blue Tank 171.94 gallons
Time(s) of watering for crops: n/a
Changes to crops: A slight growth in the tomato crop. Two existing germinations with an average size of 2.5 cm in height. Growth is observed in cucumber and lettuce crops.

– The Greenhab door opened at 9:30 am. Environmental measurements of temperature and relative humidity were taken. The GreenHab was checked and watered. The average temperature and humidity on the day was approximately 22.61 C and 11.5%. The planting of chives, cilantro, mint, parsley, rosemary begins according to the instructions received. Three pots with rainbow cherry tomato mix. Additionally, three pots of cucumber have been planted.
Support/supplies needed: None

EVA Report – November 3rd

Crew 285 EVA Report 03-11-2023
EVA # 3
Author: Andrea de la Torre
Purpose of EVA: Continue the mapping of Candor Chasma with the drone and make a 2D path to follow with the QFIELD app.
Start time: 9:00 am
End time: 11:45 am
Our mission was focused on capturing GPS coordinates throughout the journey to create a digital route that our colleagues can follow with a mobile app, leveraging the fact that one of our EVA team members was already familiar with the location. Additionally, we deployed an automated drone flight over the Candor Chasma canyon to generate an orthomosaic, a 3D model, and a DMS chart for analyzing the final exploration area for the next EVA team. Furthermore, we collected a soil sample to further our construction material project, using a different type of soil than in previous EVAs.
Destination: Candor Chasma
Coordinates (use UTM WGS 84): (520500,4251000)
· David Mateus
· Luis Diaz
· Andrea De La Torre
Road(s) and routes per MDRS Map: N Cow Dung Rd and Galileo Road.
Mode of travel: driving from the HAB north through Cow Dung Rd, turning into Galileo Road. We parked 300 meters before Compass Rock because the SOC of Perseverance was at 65% and we walked to the entrance of the canyon of Candor Chasma.

Sol Summary – November 2nd

Sol: 4
Summary Title: Talking with the Earth
Author’s name: David Mateus
Mission Status: Nominal
Sol Activity Summary: The crew started the day at 8 a.m., to have breakfast and do a communication test with Badalona Comunicacio. At 9 am, the commander and crew engineer received a filming crew from Purdue University. Due to the activities that the filming was doing around the Hab, the crew decided to stay at the science dome most of the time in order to stay in the sim.
The day was focused on the projects and preparing tomorrow’s EVAs. Also, at noon, we had an interview with Badalona Comunicacio.
Look Ahead Plan: Perform two EVA’s
Anomalies in work: None
Weather: Partly cloudy
Crew Physical Status: One of the crew members bled from their nose. You can find more details in the HSO report. The general status of the crew is nominal.
EVA: No EVA today
Reports to be filed: Sol Summary, Journalist report, Photos, Operations report, Green Hab report, Astronomy report, EVA #3 Request, EVA #4 Request, HSO Report.
Support Requested: 9V battery

Copyright © The Mars Society. All rights reserved. | Main Site